OpenStudy (anonymous):

State the domain and range of this function. y=(x+1)^2-8 I think the domain is {D=x∈R} am i right?

OpenStudy (anonymous):

Yup Domain = R ( since There is no Restriction)

OpenStudy (anonymous):

thnx does that mean the range is the same

OpenStudy (anonymous):

For Range: y =f(x) y =(x+1)^2-8 y + 8 = (x + 1)^2 sqrt ( y+8) - 1 = x nw.....here comes the restriction the sqrt of the numebr should be Real so : y + 8 > =0 y > = -8 Range = [-8 , infinity)

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

\[y=(x=1)^2-8\] \[=\sqrt{y+8}=(x-1)\] \[R={y \in R/y > -8}\] i feel like im missing a step in between last and second last equations

OpenStudy (anonymous):

it is + there (x + 1)

OpenStudy (anonymous):

oh yes sorry but im still a little confused